OpenStudy (anonymous):

Use the table to find a prediction equation to show how the number of hours worked relates to the amount of savings. Hours 4 6 10 12 15 Savings 58 82 130 154 190 a.) y=-1/2 x+ 58 b) y = 1/2x + 10 c) y=12x +58 d) y=12x+10

OpenStudy (amistre64):

well, all the options are linear; is the slope the same between all the data points?

OpenStudy (anonymous):

im not sure, that's just what the question says

OpenStudy (amistre64):

(h,s) ------- (6,82) -(4,58) -------- (2,24) ; slope = s/h = 24/2 = 12 (10,130) - ( 4, 58) -------- (6 , 72) ; slope = s/h = 72/6 = 12 so its either c or d from that experiment

OpenStudy (amistre64):

lets use (4,58) into the equation of d to check for it :) y=12x+10 58 = 12(4)+10 58 = 48 + 10 58 = 58 ; so (d) works out good

OpenStudy (anonymous):

wow thanks for the explanation